I need help with these word problems by the 14th please HELP!!!
At the beginning of a walk, Roberto and Juanita are 3.3 miles apart. If they leave at the same time and walk in the same direction, Roberto overtakes Juanita in 11 hours. If they meet in 1 hour what are their speeds?
I got 3.3x+3.3y=11, am I on the right path? and where do I go from here?
If a plane can travel 480 miles per hour with the wind and 440 miles per hour against the wind, find the speed of the plane in still air.

Answer by Alan3354(69443)

At the beginning of a walk, Roberto and Juanita are 3.3 miles apart. If they leave at the same time and walk in the same direction, Roberto overtakes Juanita in 11 hours. If they meet in 1 hour what are their speeds?
I got 3.3x+3.3y=11, am I on the right path? and where do I go from here?
r = Roberto's speed, j = Juanita's speed
r+j = 3.3
r-j = 0.3 (3.3/11)
---------- Add
2r = 3.6
r = 1.8 mph
j = 1.5 mph
------------------------
If a plane can travel 480 miles per hour with the wind and 440 miles per hour against the wind, find the speed of the plane in still air.
It's the average, (480 + 440)/2 = 460 mph
It's call airspeed, not "speed in calm air"